# A Bayesian Approach to Segmentation with Noisy Labels via Spatially Correlated Distributions

### v1

@@ 5 Conclusion
Our ELBO Computable Discrete Distribution extends beyond 2D segmentation to probabilistic models where priors are effectively represented by correlated discrete latent variables. Future work will explore its applications in 3D segmentation, volumetric data analysis, and other vision tasks that rely on structured priors incorporating spatial correlations.
@@ Acknowledgments
This research work was financially supported by the Ministry of Internal Affairs and Communications of Japan with a scheme of “Research and development of advanced technologies for a user-adaptive remote sensing data platform” (JPMI00316). We also sincerely thank Sol Cummings, Yasunori Kudo, and Katsuhiko Ishiguro for their valuable comments and suggestions on this work.
@@ A.1 KMS Matrix
The Kac–Murdock–Szegö (KMS) matrix [5, 17] is defined for a positive integer $n\in\mathbb{N}$ and a parameter $\rho\in(-1,1)$ as
@@
=$=\Bigl{\{}\prod_{i}p_{\bm{\theta}}(y^{*}_{i}|\mathbf{x})\Bigr{\}}\,\Bigl{\{}\prod_{j}p(y_{j}|y^{*}_{j},\eta_{j})\Bigr{\}}\,p(\bm{\eta}),$ (12)
@@
We introduce the variational distribution $q_{\bm{\phi}}$ as
$q_{\bm{\phi}}:=q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})=q(\bm{\eta}|\mathbf{y})\,q(\mathbf{y}^{*}|\bm{\eta},\mathbf{y})=q(\bm{\eta})\prod_{i}q(y^{*}_{i}|\eta_{i},y_{i}),$ (13)
@@
$=\log\mathbb{E}_{q_{\bm{\phi}}}\left[\frac{p_{\bm{\theta}}(\mathbf{y}|\bm{\eta},\mathbf{y}^{*},\mathbf{x})\,p_{\bm{\theta}}(\bm{\eta},\mathbf{y}^{*}|\mathbf{x})}{q_{\bm{\phi}}}\right]$ (15)
$\geq\mathbb{E}_{q_{\bm{\phi}}}\left[\log p_{\bm{\theta}}(\mathbf{y}|\bm{\eta},\mathbf{y}^{*},\mathbf{x})\right]-KL\Bigl{[}q_{\bm{\phi}}\,\|\,p_{\bm{\theta}}(\bm{\eta},\mathbf{y}^{*}|\mathbf{x})\Bigr{]}.$ (16)
@@
$\mathbb{E}_{q_{\bm{\phi}}}\left[\log p_{\bm{\theta}}(\mathbf{y}|\bm{\eta},\mathbf{y}^{*},\mathbf{x})\right]$ (17)
$=\mathbb{E}_{q_{\bm{\phi}}}\left[\log p_{\bm{\theta}}(\mathbf{y}|\bm{\eta},\mathbf{y}^{*},\mathbf{x})\right]$ (18)
$=\mathbb{E}_{q_{\bm{\phi}}}\left[\log p(\mathbf{y}|\bm{\eta},\mathbf{y}^{*})\right]$ (19)
$=\mathbb{E}_{q_{\bm{\phi}}}\left[\log\left(\prod_{i}p(y_{i}|\eta_{i},y^{*}_{i})\right)\right]$ (20)
$=\mathbb{E}_{q_{\bm{\phi}}}\left[\sum_{i}\log p(y_{i}|\eta_{i},y^{*}_{i})\right]$ (21)
$=\sum_{i}\mathbb{E}_{q_{\bm{\phi}}}\left[\log p(y_{i}|\eta_{i},y^{*}_{i})\right]$ (22)
$=\sum_{i}\mathbb{E}_{q(\eta_{i},y^{*}_{i}|\mathbf{y})}\left[\log p(y_{i}|\eta_{i},y^{*}_{i})\right]\quad(\text{marginalized by\ }(\bm{\eta}_{\backslash i},\mathbf{y}^{*}_{\backslash i}))$ (23)
@@
$=\sum_{i}\mathbb{E}_{q(y^{*}_{i}|\eta_{i},y_{i})q(\eta_{i})}\left[\log p(y_{i}|\eta_{i},y^{*}_{i})\right],\quad(y^{*}_{i}\perp\!\!\!\perp\mathbf{y}_{\backslash i}|\eta_{i},y_{i})$ (25)
where \$\backslash i$ denotes pixel indices other than $i$ on the image. We also decompose the negative KL divergence term as:
$-KL\left[q_{\bm{\phi}}\|p_{\bm{\theta}}(\bm{\eta},\mathbf{y}^{*}|\mathbf{x})\right]$ (26)
$=\mathbb{E}_{q_{\bm{\phi}}}\left[\log\frac{p_{\bm{\theta}}(\bm{\eta},\mathbf{y}^{*}|\mathbf{x})}{q_{\bm{\phi}}}\right]$ (27)
$=\mathbb{E}_{q_{\bm{\phi}}}\left[\log p_{\bm{\theta}}(\bm{\eta},\mathbf{y}^{*}|\mathbf{x})-\log q_{\bm{\phi}}\right]$ (28)
$=\mathbb{E}_{q_{\bm{\phi}}}\left[\log\underbrace{p_{\bm{\theta}}(\mathbf{y}^{*}|\bm{\eta},\mathbf{x})}_{=\ p_{\bm{\theta}}(\mathbf{y}^{*}|\mathbf{x})}\underbrace{p_{\bm{\theta}}(\bm{\eta}|\bm{\mathbf{x}})}_{=\ p(\bm{\eta})}-\log q(\mathbf{y}^{*}|\mathbf{y},\bm{\eta})\underbrace{q(\bm{\eta}|\mathbf{y})}_{=\ q(\bm{\eta})}\right]$ (29)
$=\mathbb{E}_{q_{\bm{\phi}}}\left[\log\underbrace{p_{\bm{\theta}}(\mathbf{y}^{*}|\mathbf{x})}_{=\ \prod_{i}p_{\bm{\theta}}(y_{i}^{*}|\mathbf{x})}+\log p(\bm{\eta})-\log\underbrace{q(\mathbf{y}^{*}|\mathbf{y},\bm{\eta})}_{=\ \prod_{i}\left\{q(y_{i}^{*}|y_{i},\eta_{i})\right\}}-\log q(\bm{\eta})\right]$ (30)
$=\mathbb{E}_{q_{\bm{\phi}}}\left[\sum_{i}\log p_{\bm{\theta}}(y_{i}^{*}|\mathbf{x})+\log p(\bm{\eta})-\sum_{i}\log q(y_{i}^{*}|y_{i},\eta_{i})-\log q(\bm{\eta})\right]$ (31)
$=\sum_{i}\mathbb{E}_{q(\eta_{i})q(y_{i}^{*}|\eta_{i},y_{i})}\left[\log p_{\bm{\theta}}(y_{i}^{*}|\mathbf{x})\right]+\mathbb{E}_{q(\bm{\eta})}\left[\log p(\bm{\eta})\right]$ (32)
$-\sum_{i}\mathbb{E}_{q(\eta_{i})q(y_{i}^{*}|\eta_{i},y_{i})}\left[\log q(y_{i}^{*}|y_{i},\eta_{i})\right]-\mathbb{E}_{q(\bm{\eta})}\left[\log q(\bm{\eta})\right]$ (33)
@@
$\log p_{\bm{\theta}}(\mathbf{y}|\mathbf{x})$ $\geq\sum_{i}\mathbb{E}_{q(\eta_{i})\,q(y^{*}_{i}|y_{i},\eta_{i})}\left[\log p(y_{i}|\eta_{i},y^{*}_{i})\right]$
$\quad+\sum_{i}\mathbb{E}_{q(\eta_{i})\,q(y^{*}_{i}|y_{i},\eta_{i})}\left[\log p_{\bm{\theta}}(y^{*}_{i}|\mathbf{x})\right]+\mathbb{E}_{q(\bm{\eta})}\left[\log p(\bm{\eta})\right]$
$\quad-\sum_{i}\mathbb{E}_{q(\eta_{i})\,q(y^{*}_{i}|y_{i},\eta_{i})}\left[\log q(y^{*}_{i}|y_{i},\eta_{i})\right]-\mathbb{E}_{q(\bm{\eta})}\left[\log q(\bm{\eta})\right].$ (34)

### v3

@@ 6 Conclusion
The ECCD we introduced provides a new, powerful tool for probabilistic modeling with applications extending far beyond 2D segmentation. It offers a tractable framework for problems where priors on correlated discrete variables are defined over structured domains, such as spatial grids or temporal sequences. Potential avenues for future work include extending this framework to higher-dimensional spatial problems, like 3D segmentation and volumetric data analysis. Furthermore, as the underlying KMS matrix formulation is naturally suited for modeling exponentially decaying dependencies, the ECCD holds significant promise for domains beyond computer vision, particularly for modeling temporal correlations in time-series analysis and other sequential data.
@@
Broader Impact Statement
This work introduces the ECCD, a framework designed to enhance the robustness of machine learning models against spatially correlated label noise. We believe this research has the potential for significant positive societal impact, particularly in domains where high-quality data annotation is a major bottleneck. In medical imaging, our method could lower the barrier to developing diagnostic support tools by enabling the use of larger, more diverse, and imperfectly labeled datasets, potentially accelerating research and improving healthcare accessibility. Similarly, in remote sensing, the ECCD can facilitate more accurate and large-scale analysis of satellite imagery for applications such as environmental monitoring, disaster response, and urban planning, by effectively handling noise inherent in large-scale data collection.
At the same time, we recognize potential negative repercussions that users of this technology should consider. In high-stakes applications like medical diagnosis, over-reliance on the output of any automated system, including one trained with our method, can lead to automation bias and potential misdiagnosis if not used with expert oversight. We emphasize that this technology should be used as a decision-support tool to assist qualified professionals, not to replace them. Furthermore, like many remote sensing technologies, improved segmentation capabilities could be repurposed for surveillance or other applications with negative societal consequences. We advocate for the ethical use of this technology in accordance with established legal and ethical guidelines. Finally, the performance of our method relies on the assumption of stationary, exponentially decaying correlations in the label noise. Users should be aware that in scenarios where this assumption does not hold, the model’s effectiveness may be reduced.
@@ Acknowledgments
This research work was financially supported by the Ministry of Internal Affairs and Communications of Japan with a scheme of “Research and development of advanced technologies for a user-adaptive remote sensing data platform” (JPMI00316).
@@ A.1 KMS Matrix
The Kac–Murdock–Szegö (KMS) matrix Fikioris (2018); Kac et al. (1953) is defined for a positive integer $n\in\mathbb{N}$ and a parameter $\rho\in(-1,1)$ as
@@
$=\Bigl\{\prod_{i}p_{\bm{\theta}}(y^{*}_{i}|\mathbf{x})\Bigr\}\,\Bigl\{\prod_{j}p(y_{j}|y^{*}_{j},\eta_{j})\Bigr\}\,p(\bm{\eta}),$ (23)
@@
We introduce the variational distribution $q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})$ as
$q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y}):=q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})=q(\bm{\eta}|\mathbf{y})\,q(\mathbf{y}^{*}|\bm{\eta},\mathbf{y})=q(\bm{\eta})\prod_{i}q(y^{*}_{i}|\eta_{i},y_{i}),$ (24)
@@
$=\log\mathbb{E}_{q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})}\left[\frac{p_{\bm{\theta}}(\mathbf{y}|\bm{\eta},\mathbf{y}^{*},\mathbf{x})\,p_{\bm{\theta}}(\bm{\eta},\mathbf{y}^{*}|\mathbf{x})}{q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})}\right]$ (26)
$\geq\mathbb{E}_{q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})}\left[\log p_{\bm{\theta}}(\mathbf{y}|\bm{\eta},\mathbf{y}^{*},\mathbf{x})\right]-KL\Bigl[q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})\,\|\,p_{\bm{\theta}}(\bm{\eta},\mathbf{y}^{*}|\mathbf{x})\Bigr].$ (27)
@@
$\mathbb{E}_{q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})}\left[\log p_{\bm{\theta}}(\mathbf{y}|\bm{\eta},\mathbf{y}^{*},\mathbf{x})\right]$ (28)
$=\mathbb{E}_{q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})}\left[\log p_{\bm{\theta}}(\mathbf{y}|\bm{\eta},\mathbf{y}^{*},\mathbf{x})\right]$ (29)
$=\mathbb{E}_{q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})}\left[\log p(\mathbf{y}|\bm{\eta},\mathbf{y}^{*})\right]$ (30)
$=\mathbb{E}_{q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})}\left[\log\left(\prod_{i}p(y_{i}|\eta_{i},y^{*}_{i})\right)\right]$ (31)
$=\mathbb{E}_{q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})}\left[\sum_{i}\log p(y_{i}|\eta_{i},y^{*}_{i})\right]$ (32)
$=\sum_{i}\mathbb{E}_{q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})}\left[\log p(y_{i}|\eta_{i},y^{*}_{i})\right]$ (33)
[…]
@@
[…]
$-KL\left[q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})\|p_{\bm{\theta}}(\bm{\eta},\mathbf{y}^{*}|\mathbf{x})\right]$ (37)
$=\mathbb{E}_{q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})}\left[\log\frac{p_{\bm{\theta}}(\bm{\eta},\mathbf{y}^{*}|\mathbf{x})}{q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})}\right]$ (38)
$=\mathbb{E}_{q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})}\left[\log p_{\bm{\theta}}(\bm{\eta},\mathbf{y}^{*}|\mathbf{x})-\log q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})\right]$ (39)
$=\mathbb{E}_{q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})}\left[\log\underbrace{p_{\bm{\theta}}(\mathbf{y}^{*}|\bm{\eta},\mathbf{x})}_{=\ p_{\bm{\theta}}(\mathbf{y}^{*}|\mathbf{x})}\underbrace{p_{\bm{\theta}}(\bm{\eta}|\bm{\mathbf{x}})}_{=\ p(\bm{\eta})}-\log q(\mathbf{y}^{*}|\mathbf{y},\bm{\eta})\underbrace{q(\bm{\eta}|\mathbf{y})}_{=\ q(\bm{\eta})}\right]$ (40)
$=\mathbb{E}_{q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})}\left[\log\underbrace{p_{\bm{\theta}}(\mathbf{y}^{*}|\mathbf{x})}_{=\ \prod_{i}p_{\bm{\theta}}(y_{i}^{*}|\mathbf{x})}+\log p(\bm{\eta})-\log\underbrace{q(\mathbf{y}^{*}|\mathbf{y},\bm{\eta})}_{=\ \prod_{i}\left\{q(y_{i}^{*}|y_{i},\eta_{i})\right\}}-\log q(\bm{\eta})\right]$ (41)
$=\mathbb{E}_{q(\bm{\eta},\mathbf{y}^{*}|\mathbf{y})}\left[\sum_{i}\log p_{\bm{\theta}}(y_{i}^{*}|\mathbf{x})+\log p(\bm{\eta})-\sum_{i}\log q(y_{i}^{*}|y_{i},\eta_{i})-\log q(\bm{\eta})\right]$ (42)
$=-\mathbb{E}_{q(\bm{\eta})}\left[\log q(\bm{\eta})\right]+\mathbb{E}_{q(\bm{\eta})}\left[\log p(\bm{\eta})\right]-\sum_{i}\mathbb{E}_{q(\eta_{i})q(y_{i}^{*}|\eta_{i},y_{i})}\left[\log q(y_{i}^{*}|y_{i},\eta_{i})\right]$ (43)
$\quad+\sum_{i}\mathbb{E}_{q(\eta_{i})q(y_{i}^{*}|\eta_{i},y_{i})}\left[\log p_{\bm{\theta}}(y_{i}^{*}|\mathbf{x})\right]$ (44)
@@
$\log p_{\bm{\theta}}(\mathbf{y}|\mathbf{x})$ $\geq\sum_{i}\mathbb{E}_{q(\eta_{i})\,q(y^{*}_{i}|y_{i},\eta_{i})}\left[\log p(y_{i}|\eta_{i},y^{*}_{i})\right]-\mathbb{E}_{q(\bm{\eta})}\left[\log q(\bm{\eta})\right]+\mathbb{E}_{q(\bm{\eta})}\left[\log p(\bm{\eta})\right]$
$\quad-\sum_{i}\mathbb{E}_{q(\eta_{i})\,q(y^{*}_{i}|y_{i},\eta_{i})}\left[\log q(y^{*}_{i}|y_{i},\eta_{i})\right]+\sum_{i}\mathbb{E}_{q(\eta_{i})\,q(y^{*}_{i}|y_{i},\eta_{i})}\left[\log p_{\bm{\theta}}(y^{*}_{i}|\mathbf{x})\right].$ (45)